How do you convert the vertex form y=4(3x-1)^2+8 back into standard form?
Found 2 solutions by greenestamps, josgarithmetic:
Answer by greenestamps(13195)



Expand the expression on the right and combine like terms if there are any.

(3x-1)^2 = 9x^2-6x+1
4(9x^2-6x+1)+8 = (36x^2-24x+4)+8 = 36x^2-24x+12

ANSWER: y=36x^2-24x+12


Answer by josgarithmetic(39613)

y equals, the other side is to be fully multiplied and then simplified in decreasing order of the single variable x.